OpenStudy (anonymous):

Indefinite integral of x/(1+sqrt(x))?

OpenStudy (anonymous):

2sqrt(x) - 2ln(1 + sqrt(x)) + C u = 1 + sqrt(x)

OpenStudy (anonymous):

Jamesm, how did you get that? I didn't get anything close to that lol.

OpenStudy (anonymous):

what did you get? try differentiating it to see if it's correct.

OpenStudy (anonymous):

but i got 1/2sqrt(x) if i differentiate mr jamesm

OpenStudy (anonymous):

I got: \[\frac{2}{3}x^{\frac{3}{2}}-x+2\sqrt{x}-2\ln|\sqrt{x}+1|+C\]

OpenStudy (anonymous):

you are correct. i integrated 1/(sqrt(x) + 1) instead. didn't read it correctly. :)

OpenStudy (anonymous):

Oops, my fault, after jamesm previous effort I just glanced at his answer, saw the ln and his substitution and assumed it was right. So sorry.